You invested $30,000 in a mutual fund at the beginning of the year when the NAV was $55.72. At the end of the year, the fund paid $0.28 in short-term distributions and $0.45 in long-term distributions. If the NAV of the fund at the end of the year was $60.84, what was your return for the year?
Answer #1

no of unit purchased = 30000/55.72

return for the year in dollars = number of unit purchased * (short term distributions + long term distributions + NAV at the end - NAV at the time of purchase)

= 30000/55.72 * (0.28+0.45+60.84-55.72)

= 3149.68

Return for the year = 3149.68/30000

= 10.50%